Determine the angle 0 between the sides of the tri
Answer #1

We have the following points: A(0, 1, 1), B(0, 3, 4) and C(3, 5, 0)

Now,

vector AB = (0 - 0) i + (3 - 1) j + (4 - 1) k = 2 j + 3 k

vector AC = (3 - 0) i + (5 - 1) j + (0 - 1) k = 3 i + 4 j - k

Using the dot product of vectors AB and AC, we get,

cosθ = AB.AC / (|AB| |AC|) = (2 j + 3 k).(3 i + 4 j - k) / [(22 + 32)1/2 (32 + 42 + (-1)2)1/2] = 5 / (13 * 21/2)

=> θ = cos-1[5 / (13 * 21/2)] = 74.2o